# SOSecure: Safer Code Generation with RAG and StackOverflow Discussions

### v1

@@ 4.1. Baseline Systems under test
4.2. Benchmark datasets
SALLM contains 100 prompts, available in both text and code formats, along with the corresponding generated code snippets covering 45 vulnerability types (CWEs). Each snippet is mapped to a CWE. We select samples that include default CodeQL QL files, resulting in a final set of 74 samples used for analysis.
LMSYS-Chat-1M consists of 1 million samples, of which 43,269 contain Python code. Among these, 31,008 samples are single-round user conversations with a single code block. To curate a high-quality dataset containing genuine vulnerabilities, we apply a two-step filtration process. First, we run the samples through two static analyzers (Bandit (PyCQA, 2024) and CodeQL) and retain only those flagged as vulnerable by both. This results in 2,809 samples. We further select samples associated with CWEs that include default CodeQL QL files, yielding a final set of 240 Python samples for analysis.
4.3. Evaluation Metrics
To determine the presence or absence of security issues in the generated code, we use CodeQL (GitHub, 2025) for evaluation. CodeQL is a static analysis tool designed to automatically detect vulnerabilities by executing QL queries on a database generated from the source code. We also study two programming languages, Python and C. This served as an additional motivating factor for using CodeQL, as it allows for a uniform analysis, enabling the use of a single tool to assess both languages. We compute the following security metrics to evaluate SOSecure:
Fix Rate (FR). % of security issues that were fixed by the system. A higher FR indicates better security performance.
Intro Rate (IR). % of new security issues introduced after code generation. A lower IR reflects fewer new vulnerabilities.
secure@k, vulnerable@k (Siddiq et al., 2024b) measures the security of generated code. vulnerable@k measures the probability that at least one code snippet out of k generated samples is vulnerable. For this metric, a lower score indicates better performance of the system. The secure@k metric measures the probability that all code snippets out of k samples are vulnerability-free. For this metric, a higher score indicates better performance of the system. We use both these metrics to compare the performance of SOSecure and SALLM.
5. Results
| Temperature | Metric | SOSecure (%) | SALLM (%) |
| 0.0 | secure@1 | 89.19 | 51.35 |
| vulnerable@1 | 10.81 | 48.65 | |
| 0.2 | secure@1 | 85.14 | 50.0 |
| vulnerable@1 | 14.86 | 50.0 | |
| 0.4 | secure@1 | 83.78 | 51.35 |
| vulnerable@1 | 16.22 | 48.65 | |
| 0.6 | secure@1 | 79.73 | 51.35 |
| vulnerable@1 | 20.27 | 48.65 | |
| 0.8 | secure@1 | 85.14 | 52.70 |
| vulnerable@1 | 14.86 | 47.30 | |
| 1.0 | secure@1 | 75.68 | 52.70 |
| vulnerable@1 | 24.32 | 47.30 |
| SOSecure | GPT4 | GPT4+CWE | |||||
| CWE | Sev | F/T | P% | F/T | P% | F/T | P% |
| 078 | 9.8 | 6/6 | 100 | 2/6 | 33.3 | 3/6 | 50 |
| 190 | 8.6 | 2/2 | 100 | 2/2 | 100 | 2/2 | 100 |
| 022 | 7.5 | 0/3 | 0 | 0/3 | 0 | 0/3 | 0 |
| Avg | 66.7 | 44.4 | 50 | ||||
| System | FR (%) | IR (%) | NCR (%) |
| SOSecure | 73.3 | 0 | 72.5 |
| GPT4 | 53.3 | 0 | 80 |
| GPT4+CWE | 60 | 0 | 77.5 |
| GPT4+CWE+ | 53.3 | 0 | 80 |
5.1. Effectiveness of SOSecure
Table 1 presents the comparison of SOSecure with baseline approaches across three benchmark datasets. The results demonstrate that SOSecure consistently outperforms standard LLM approaches in terms of security vulnerability mitigation. On the SALLM dataset, SOSecure achieves a Fix Rate (FR) of 71.7%, significantly higher than both GPT4 (49.06%) and GPT+CWE (58.49%). Similarly SOSecure achieves a 91.3% Fix Rate compared to 56.52% for GPT4 and 69.57% for GPT+CWE. The most substantial performance gain is observed on the LMSys dataset, where SOSecure achieves a remarkable 96.67% Fix Rate, compared to only 37.50% for GPT4 and 45.83% for GPT+CWE. Importantly, while improving security, SOSecure does not introduce new vulnerabilities, maintaining a 0% Introduced vulnerabilities Rate (IR) across all datasets. In contrast, GPT+CWE introduces new vulnerabilities in 7.69% of cases on the LLMSecEval dataset. The No Change Rate (NCR) is also consistently lower for SOSecure (48.65%, 57.14%, and 3.33% across the three datasets) compared to both baseline approaches.
Table 3 provides a breakdown of performance by CWE type across the three datasets. The results reveal that SOSecure demonstrates varying effectiveness depending on the vulnerability type, with particularly strong performance on high-severity vulnerabilities. In the SALLM dataset, SOSecure achieves perfect precision (100%) for CWE-918 (Server-Side Request Forgery) and CWE-089 (SQL Injection), both high-severity vulnerabilities with CVSS scores of 9.1 and 8.8, respectively. For CWE-094 (Code Injection, severity 9.3), SOSecure achieves 77.78% precision, outperforming GPT4 (55.56%) while matching GPT+CWE. SOSecure also demonstrates superior performance for CWE-078 (OS Command Injection, severity 6.3), achieving 90% precision compared to 80% for both baseline approaches. Similar patterns are observed in the LMSys dataset, where SOSecure achieves 100% precision for CWE-094 and strong performance across other vulnerability types. In the LLMSecEval dataset, SOSecure achieves perfect precision for five out of seven CWE types, including the high-severity vulnerabilities CWE-502 (Deserialization of Untrusted Data, severity 9.8) and CWE-798 (Use of Hard-coded Credentials, severity 9.8). Notably, SOSecure demonstrates consistent improvement across vulnerability types of varying severity levels, indicating that the SO discussions provide valuable security insights across a broad spectrum of security concerns.
5.2. Performance By Vulnerability Type
In the SALLM dataset, SOSecure achieves perfect precision (100%) for CWE-918 (Server-Side Request Forgery) and CWE-089 (SQL Injection), which are critical (9.1) and high (8.8) severity vulnerabilities, respectively. For CWE-094 (Code Injection, severity 9.3), a critical vulnerability, SOSecure achieves 77.78% precision, outperforming GPT4 (55.56%) while matching GPT+CWE. SOSecure also demonstrates superior performance for CWE-078 (OS Command Injection, severity 6.3), a medium severity vulnerability, achieving 90% precision compared to 80% for both baseline approaches.
5.3. Impact of Number of Retrieved Candidates
Figure 5 illustrates the effect of varying the number of neighbors (k) added as context on the average precision of SOSecure for the LMSys data. We observe that performance initially increases with the number of neighbors, peaking at around k=5 with approximately 94% average precision. Beyond this point, adding more neighbors leads to a decline in performance, suggesting that excessive context may introduce noise or conflicting information. Too few discussions may not provide sufficient security insights, while too many may dilute the focus on the specific vulnerability being addressed.
@@ 5.4. Performance Across Languages
To evaluate the language generalizability of our approach, we tested SOSecure on both Python and C code samples from the LLMSecEval dataset. Results are presented in Table 4. SOSecure maintained strong performance across both languages, achieving a Fix Rate (FR) of 73.33% on C code samples compared to 53.33% for GPT4 and 60% for GPT+CWE. Similar to our findings with Python, SOSecure maintained a 0% Introduced vulnerabilities Rate (IR) on C code. When analyzed by CWE type for C code, SOSecure achieved 100% precision for CWE-078 (OS Command Injection) and CWE-190 (Integer Overflow), demonstrating that our approach generalizes well across different programming languages. However, SOSecure showed limitations with CWE-022 (Path Traversal) in C code, similar to our findings with Python. These results suggest that SOSecure’s approach of leveraging Stack Overflow discussions as context is effective across multiple programming languages.
@@ 6.1.1. Applications of Community Security Insights
Although the input code already used pickle, SOSecure correctly recalled that pickle also present security risks and transformed the code to use safer JSON serialization:
@@ 6.1.2. Patterns in Effective Community Knowledge Transfer
Across the analyzed examples, several patterns emerged in how SOSecure effectively leverages community knowledge (when backed by a capable LLM). When SO comments explicitly mention security concerns (e.g., ”Don’t use shell=True”), SOSecure directly applies these insights in its fixes. Even when comments don’t explicitly mention a vulnerability type, they often provide contextual information about proper framework usage that SOSecure combines with security best practices. SOSecure doesn’t blindly apply all suggestions from SO, but critically evaluates them for security implications, as seen in the pickle/eval example. Many effective fixes leverage community discussions about framework-specific practices, such as Flask’s debug mode, demonstrating how framework expertise contributes to security. These patterns highlight the various mechanisms through which community knowledge enhances the security of LLM-generated code.
@@ 6.2. Qualitative Analysis of Failure Cases
The SO discussions didn’t mention this security concern, providing no guidance on proper host key validation for SSH connections. This failure demonstrates that even when SOSecure can improve some aspects of security (credential handling), it may miss other critical vulnerabilities when the community discussions don’t address them.
@@ 6.2.1. Patterns in Failure Cases
Analyzing these and other failure cases reveals several recurring patterns in situations where SOSecure is unable to fully remediate security vulnerabilities. When community discussions reference outdated security practices that were once considered secure but are now vulnerable (like TLSv1), SOSecure cannot always discern that these recommendations need to be updated to current standards. SOSecure struggles when secure implementations require knowledge that isn’t explicitly stated in the retrieved SO discussions, which is particularly challenging for domain-specific security practices. In many failure cases, SOSecure makes partial improvements to security, such as replacing hardcoded credentials or improving error handling, but misses deeper architectural or protocol-level vulnerabilities that require specialized security expertise. When security decisions involve trade-offs between usability and security that are context-dependent, SOSecure may not have sufficient information to make the optimal decision. We also find that retrieval sometimes returns discussions that focus on debugging rather than security, leading to incomplete fixes. Future work may be able to address these limitations by sampling multiple, complementary discussions to be used in the prompt rather than just the top-k neighbors. For many patterns, it may even be beneficial to produce aggregate all security insights from SO discussions into regularly updated knowledge bases to produce shorter and more effective rewrite prompts.
@@ 6.3. Code Similarity Analysis
Since most of the evaluation datasets do not have test cases, we rely on extensive manual annotation and careful prompting to ensure that SOSecure does not simply remove risky portions of code to address the vulnerability. We did not find evidence of such behavior. We also quantify the average extent of modifications made to the code by SOSecure by calculating the difference between the original and fixed code for the LMSys dataset. We find that successful vulnerability repairs maintained an average similarity of 0.60 (using Python’s difflib library, on a scale from 0 to 1), indicating that SOSecure is making targeted security modifications rather than extensive rewrites. We are therefore reasonably confident that it preserves the developer’s original intent while addressing specific security concerns, when backed by models at least as capable as GPT4 (used in this work).
@@ 6.4. The Importance of Community Forums
The performance of SOSecure across all three datasets validates our core hypothesis that community-driven security insights can meaningfully improve LLM code generation. This effectiveness stems from the key advantages of SO discussions over LLMs: they naturally capture evolving security practices and adapt quickly to newly discovered vulnerabilities. The significant improvement in fix rates for high-severity vulnerabilities (e.g., CWE-094, CWE-502) suggests that SO discussions provide valuable contextual information about security implications specific to particular libraries or API usage patterns, with this contextual awareness evident in cases like CWE-078 (OS Command Injection), where community comments explicitly warn against unsafe practices. Additionally, the collective wisdom captured in SO discussions represents insights from security experts, library maintainers, and experienced developers across various domains, with security knowledge being highly domain-specific and vulnerabilities often tied to particular frameworks, libraries, or implementation contexts. This diversity of expertise contributes to SOSecure’s ability to address a wide range of vulnerability types effectively.
An important implication of our work relates to the evolving relationship between community knowledge platforms like SO and AI code generation tools. As developers increasingly rely on LLMs for code generation, forums like SO have seen a steep decrease in usage. There is a risk that fewer new discussions will be created on community platforms, not to mention that more and more of these discussions will be AI-authored, potentially diminishing this valuable knowledge resource. As shown, this would harm AI-based methods as well. SOSecure demonstrates the continued value of community discussions as a complement to LLM capabilities.
@@ 6.5. Limitations
CodeQL accuracy constraints: Our evaluation relies on CodeQL for vulnerability detection, which may have false positives and negatives. Additionally, our analysis is limited to CWEs with existing default CodeQL queries, potentially missing other important vulnerability types. Future work could explore complementary analysis techniques to provide more comprehensive security evaluation.
Retrieval strategy: SOSecure’s effectiveness depends on the availability and quality of security-related discussions in SO. For newer or niche technologies with limited community discourse, the approach may be less effective. Additionally, we use BM25 as the retrieval mechanism. SOSecure could also benefit from developing more sophisticated filtering mechanisms for context selection.
@@ 7. Conclusion
This paper introduces SOSecure, a retrieval-augmented approach for improving security in LLM-generated code using community knowledge from Stack Overflow. SOSecure constructs a security-oriented knowledge base from posts and comments containing explicit security warnings, retrieves relevant discussions, and incorporates them as context during code revision. Our evaluation across three datasets and two languages demonstrates SOSecure’s effectiveness in mitigating security vulnerabilities. SOSecure does not require retraining or specialized fine-tuning, allowing seamless integration into existing LLM deployments with minimal overhead. While the primary evaluation focused on Python, the results from the C language dataset (Table 4) show that SOSecure can generalize across programming languages. Additionally, as security discussions evolve on platforms like Stack Overflow, SOSecure’s knowledge base can be continuously updated, ensuring ongoing improvements in security without the need for retraining the model.

### v2

@@ 6.4. Performance Across Languages
To evaluate the language generalizability of our approach, we tested SOSecure on both Python and C code samples from the LLMSecEval dataset. Results are presented in Table 3. SOSecure maintained strong performance across both languages, achieving a Fix Rate (FR) of 73.33% on C code samples compared to 53.33% for GPT4 and 60% for GPT4+CWE. Similar to our findings with Python, SOSecure maintained a 0% Introduced vulnerabilities Rate (IR) on C code. When analyzed by CWE type for C code, SOSecure achieved 100% precision for CWE-078 (OS Command Injection) and CWE-190 (Integer Overflow), demonstrating that our approach generalizes well across different programming languages. However, SOSecure showed limitations with CWE-022 (Path Traversal) in C code, similar to our findings with Python. These results suggest that SOSecure’s approach of leveraging SO discussions as context is effective across multiple programming languages.
@@ 7.1.1. Applications of Community Security Insights
Although the input code already used pickle, SOSecure correctly recalled that pickle also presents security risks and transformed the code to use safer JSON serialization:
@@ 7.1.2. Patterns in Effective Community Knowledge Transfer
Across the analyzed examples, several patterns emerged in how SOSecure effectively leverages community knowledge (when backed by a capable LLM). When SO comments explicitly mention security concerns (e.g., ”Don’t use shell=True”), SOSecure directly applies these insights in its fixes. Even when comments don’t explicitly mention a vulnerability type, they often provide contextual information about proper framework usage that SOSecure combines with security best practices. SOSecure does not blindly apply all suggestions from SO, but critically evaluates them for security implications, as seen in the pickle/eval example. Many effective fixes leverage community discussions about framework-specific practices, such as Flask’s debug mode, demonstrating how framework expertise contributes to security. These patterns highlight the various mechanisms through which community knowledge enhances the security of LLM-generated code.
@@ 7.2. Qualitative Analysis of Failure Cases
The SO discussions did not mention this security concern, providing no guidance on proper host key validation for SSH connections. This failure demonstrates that even when SOSecure can improve some aspects of security (credential handling), it may miss other critical vulnerabilities when community discussions do not address them.
@@ 7.2.1. Patterns in Failure Cases
Analyzing these and other failure cases reveals several recurring patterns in situations where SOSecure is unable to fully remediate security vulnerabilities. When community discussions refer to outdated security practices that were once considered secure but are now vulnerable (such as TLSv1), SOSecure cannot always discern that these recommendations need to be updated to current standards. SOSecure struggles when secure implementations require knowledge that isn’t explicitly stated in the retrieved SO discussions, which is particularly challenging for domain-specific security practices. In many failure cases, SOSecure makes partial improvements to security, such as replacing hardcoded credentials or improving error handling, but misses deeper architectural or protocol-level vulnerabilities that require specialized security expertise. When security decisions involve trade-offs between usability and security that are context-dependent, SOSecure may not have sufficient information to make the optimal decision. We also find that retrieval sometimes returns discussions that focus on debugging rather than security, leading to incomplete fixes. Future work may be able to address these limitations by sampling multiple, complementary discussions to be used in the prompt rather than just the top-k neighbors. For many patterns, it may even be beneficial to aggregate all security insights from SO discussions into regularly updated knowledge bases to produce shorter and more effective rewrite prompts.
@@ 7.3. Code Similarity Analysis
Since most of the evaluation datasets do not have test cases, we rely on extensive manual annotation and careful prompting to ensure that SOSecure does not simply remove risky portions of code to address the vulnerability. We did not find evidence of this behavior. We also quantify the average extent of modifications made to the code by SOSecure by calculating the difference between the original and fixed code for the LMSys dataset. We find that successful vulnerability repairs maintained an average similarity of 0.60 (using Python’s difflib library, on a scale from 0 to 1), indicating that SOSecure is making targeted security modifications rather than extensive rewrites. We are therefore reasonably confident that it preserves the developer’s original intent while addressing specific security concerns, when backed by models at least as capable as GPT4 (used in this work).
@@ 7.4. The Importance of Community Forums
The performance of SOSecure across the three data sets validates our core hypothesis that community-driven security insights can meaningfully improve LLM code generation. This effectiveness stems from the key advantages of SO discussions over LLMs: they naturally capture evolving security practices and adapt quickly to newly discovered vulnerabilities. The significant improvement in fix rates for high-severity vulnerabilities (e.g., CWE-094, CWE-502) suggests that SO discussions provide valuable contextual information about security implications specific to particular libraries or API usage patterns, with this contextual awareness evident in cases like CWE-078 (OS Command Injection), where community comments explicitly warn against unsafe practices. Additionally, the collective wisdom captured in SO discussions represents insights from security experts, library maintainers, and experienced developers across various domains, with security knowledge being highly domain specific and vulnerabilities often tied to particular frameworks, libraries, or implementation contexts. This diversity of expertise contributes to SOSecure’s ability to address a wide range of vulnerability types effectively.
An important implication of our work is related to the evolving relationship between community knowledge platforms such as SO and AI code generation tools. As developers increasingly rely on LLMs for code generation, forums like SO have seen a steep decrease in usage. There is a risk that fewer new discussions will be created on community platforms, not to mention that more and more of these discussions will be AI-authored, potentially diminishing this valuable knowledge resource. As shown, this would also harm AI-based methods. SOSecure demonstrates the continued value of community discussions as a complement to LLM capabilities.
@@ 7.5. Limitations
CodeQL accuracy constraints. Our evaluation relies on CodeQL for vulnerability detection, which can yield false positives and false negatives. The analysis is further limited to CWEs covered by CodeQL’s default queries, leaving out other potentially important vulnerability classes. Consequently, our results capture improvements on vulnerabilities detectable by CodeQL, while categories such as logic flaws, side channels, or misconfigurations remain outside of its scope. Incorporating complementary analysis tools and broader evaluation criteria is a promising direction for future work.
Retrieval strategy: SOSecure’s effectiveness depends on the availability and quality of security-related discussions in SO. For newer or niche technologies with limited community discourse, the approach may be less effective.
Although we experimented with dense retrieval methods (e.g., FAISS with cosine similarity), BM25 consistently performed better in our setting, particularly for queries involving API names and error messages. We therefore adopt BM25 in SOSecure. Future work may revisit semantic or hybrid retrieval approaches that combine lexical precision with semantic recall.
@@ 8. Conclusion
This paper introduces SOSecure, a retrieval-augmented approach for improving security in LLM-generated code using community knowledge from Stack Overflow. SOSecure constructs a security-oriented knowledge base from posts and comments containing explicit security warnings, retrieves relevant discussions, and incorporates them as context during code revision. Our evaluation across three datasets and two languages demonstrates the effectiveness of SOSecure in mitigating security vulnerabilities. SOSecure does not require retraining or specialized fine-tuning, allowing seamless integration into existing LLM deployments with minimal overhead. Although the primary evaluation focused on Python, the results from the C language dataset (Table 3) show that SOSecure can generalize between programming languages. Additionally, as security discussions evolve on platforms like SO, SOSecure’s knowledge base can be continuously updated, ensuring ongoing improvements in security without the need for retraining the model.
Data Availability